# EquiTabPFN: A Target-Permutation Equivariant Prior Fitted Networks

### v1

@@ 7 Conclusion
In this paper, we discussed the property of target equivariance and its importance for foundational tabular models. We proved that having a non-equivariant model provably worsens the pre-training objective with an incompressible term and proposed an architecture that is equivariant to target permutation. We showed that our new architecture performs better than TabPFN, Mothernet and traditional tabular baselines on artificial and real-world datasets. We hope this work can be useful to interpret tabular foundational models as deep kernel machines and offer new insights to understand the performance of tabular foundational models.
We will release our code for training and evaluating our model, enabling reproducibility of our results.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Proofs
Appendix B Binary classification decision boundary
@@
Appendix C Additional experiment results
Figure 6: Boxplot of normalized AUC for all methods on the 30 classification datasets of Hollmann et al. (2023).
In Figure 6, we show the boxplot of normalized AUC in all datasets. The AUC is min-max normalized with the minimum and maximum scores observed on each dataset.
When computing critical diagram in Figure 4, we use Autorank implementation (Herbold, 2020) with default hyperparameters which corresponds to a significance level of $0.05$.
Appendix D Experimental details
did name p n q
11 balance-scale 5 625 3
14 mfeat-fourier 77 2000 10
15 breast-w 10 699 2
16 mfeat-karhunen 65 2000 10
18 mfeat-morphological 7 2000 10
22 mfeat-zernike 48 2000 10
23 cmc 10 1473 3
29 credit-approval 16 690 2
31 credit-g 21 1000 2
37 diabetes 9 768 2
50 tic-tac-toe 10 958 2
54 vehicle 19 846 4
188 eucalyptus 20 736 5
458 analcatdata_authorship 71 841 4
469 analcatdata_dmft 5 797 6
1049 pc4 38 1458 2
1050 pc3 38 1563 2
1063 kc2 22 522 2
1068 pc1 22 1109 2
1462 banknote-authentication 5 1372 2
1464 blood-transfusion-… 5 748 2
1480 ilpd 11 583 2
1494 qsar-biodeg 42 1055 2
1510 wdbc 31 569 2
6332 cylinder-bands 40 540 2
23381 dresses-sales 13 500 2
40966 MiceProtein 82 1080 8
40975 car 7 1728 4
40982 steel-plates-fault 28 1941 7
40994 climate-model-… 21 540 2
Table 3: Test dataset names and properties, taken from Hollmann et al. (2023). Here did is the OpenML Dataset ID, p the number of covariates, n the number of samples, and q the number of classes in each dataset.
did name p n q
6 letter 16 20000 26
1515 micro-mass 1300 571 20
41084 UMIST_Faces_Cropped 10304 575 20
42 soybean 35 683 19
40971 collins 23 1000 30
300 isolet 617 7797 26
1481 kr-vs-k 6 28056 18
184 kropt 6 28056 18
40499 texture 40 5500 11
Table 4: Names and properties of the datasets used for evaluating EquiTabPFN on unseen number of classes. The datasets are extracted from the OpenML-CC18 benchmark suite (Bischl et al., 2021) and have a number of classes greater than $10$ (ranging from $11$ to $30$). Here did is the OpenML Dataset ID, p the number of covariates, n the number of samples, and q the number of classes in each dataset.

### v2

@@ 7 Conclusion
Summary. In this paper, we introduced EquiTabPFN, an architecture for PFNs that is equivariant to target permutations and enables in-context learning on datasets with arbitrary class counts. We established optimality of equivariant architectures for foundational tabular models and proved that non-equivariant models worsens the pre-training objective with an incompressible error term. Finally, we empirically showed the benefits of EquiTabPFN both in terms of classification performance and inference runtime on synthetic and real-world datasets. We hope this work enables future developments on prior-fitted networks that incorporate such fundamental symmetry of tabular data.
Limitations. EquiTabPFN requires a quadratic extra-cost with the number of target dimensions to perform self-attention. While many tabular problems have a relatively small number of target dimensions, this may become a problem for very large number of dimension and future work could consider efficient self-attention to address this issue.
Code for training and evaluating our model, will be released to enable reproducibility of our results.
Acknowledgments
This work was supported by the ANR project BONSAI (grant ANR-23-CE23-0012-01) and was granted access to the HPC resources of IDRIS under the allocation [AD011015767] made by GENCI.
@@ Appendix A Proofs
Appendix B Additional Experimental details
Training procedure. We use a similar training protocol as in Hollmann et al. [9], in which the model is trained on classification datasets generated according to their proposed artificial dataset prior. In this protocol, each dataset has a fixed size of $1024$ and is split into training and test uniformly at random. The maximum number of classes is fixed to $10$, while the maximum dimension of the covariate vector is fixed to $100$. Following Müller et al. [17], we represent the target $y$ as a one-hot encoding vector whose dimension is the number of classes in the dataset. Moreover, we employ the exact same strategy for handling missing values in the covariates. Training is performed using 153600 batches of 72 synthetically generated datasets each, which means the model was exposed to $\sim$11M artificial datasets during pre-training, a similar order of magnitude of datasets used for pre-training TabPFN by Hollmann et al. [9]. The total training time of the network lasts approximately 4 days on a single A100 GPU with 80GB of GPU memory. The resulting network is then used for all our evaluations without altering its parameters.
We used the Adam optimizer [12] with initial learning rate of $0.0001$ and linear-warmup scheduler for the first $10$ epochs followed by cosine annealing [13] as in Hollmann et al. [9].
Architecture details. We use an EquiTabPFN network with $12$ self-attention layers alternating between both type of attention introduced in Section 4: 6 blocks $\text{SelfAtt}_{c}$ and 6 blocks $\text{SelfAtt}_{b}$. Each self-attention layer consists of a multi-head attention blocks with $4$ heads, embeddings of dimension $512$, and hidden layers of dimension $1024$. This choice ensures a fair comparison with the models used in Hollmann et al. [9], Müller et al. [17], since the number of parameters ($25.17M$) are of the same order when counting them as proposed in Kaplan et al. [11].
Datasets. For the datasets with less than $10$ classes, we collect the ones with less than $3000$ samples and $100$ features and retain the datasets that contain the same number of classes accross all folds and splits. For the datasets with more than $10$ classes, we filter in addition the ones able to run inference on an 80GB A100 GPU. The datasets obtained are given in Table 1 and Table 3.
taskId name Classes Features Samples
3 kr-vs-kp 2 36 2556
4 labor 2 16 45
9 autos 6 25 163
11 balance-scale 3 4 499
14 mfeat-fourier 10 76 1600
15 breast-w 2 9 559
16 mfeat-karhunen 10 64 1600
18 mfeat-morphological 10 6 1600
22 mfeat-zernike 10 47 1600
23 cmc 3 9 1177
25 colic 2 26 294
27 colic 2 22 294
29 credit-approval 2 15 552
31 credit-g 2 20 800
35 dermatology 6 34 292
37 diabetes 2 8 614
39 sonar 2 60 166
40 glass 6 9 170
45 splice 3 60 2552
47 tae 3 5 120
48 heart-c 2 13 241
49 tic-tac-toe 2 9 766
50 heart-h 2 13 234
53 vehicle 4 18 676
54 hepatitis 2 19 123
59 iris 3 4 120
2079 eucalyptus 5 19 588
2867 anneal 5 38 718
3512 synthetic-control 6 60 480
3540 analcatdata-boxing1 2 3 96
3543 irish 2 5 400
3549 analcatdata-authorship 4 70 672
3560 analcatdata-dmft 6 4 637
3561 profb 2 9 536
3602 visualizing-environmental 2 3 88
3620 fri-c0-100-5 2 5 80
3647 rabe-266 2 2 96
3731 visualizing-livestock 2 2 104
taskId name Classes Features Samples
3739 analcatdata-chlamydia 2 3 80
3748 transplant 2 3 104
3779 fri-c3-100-5 2 5 80
3797 socmob 2 5 924
3902 pc4 2 37 1166
3903 pc3 2 37 1249
3913 kc2 2 21 416
3917 kc1 2 21 1687
3918 pc1 2 21 887
9946 wdbc 2 30 455
9957 qsar-biodeg 2 41 843
9971 ilpd 2 10 465
9978 ozone-level-8hr 2 72 2026
9979 cardiotocography 10 35 1700
9984 fertility 2 9 80
10089 acute-inflammations 2 6 96
10093 banknote-authentication 2 4 1096
10101 blood-transfusion-service-center 2 4 598
14954 cylinder-bands 2 37 432
14967 cjs 6 33 2236
125920 dresses-sales 2 12 400
125921 LED-display-domain-7digit 10 7 400
145793 yeast 4 8 1015
145799 breast-cancer 2 9 228
145836 blood-transfusion-service-center 2 4 598
145847 hill-valley 2 100 968
145984 ionosphere 2 34 280
146024 lung-cancer 3 56 24
146063 hayes-roth 3 4 128
146065 monks-problems-2 2 6 480
146192 car-evaluation 4 21 1382
146210 postoperative-patient-data 2 8 70
146800 MiceProtein 8 77 864
146817 steel-plates-fault 7 27 1552
146818 Australian 2 14 552
146819 climate-model-simulation-crashes 2 18 432
146821 car 4 6 1382
146822 segment 7 16 1848
Table 1: List of the $76$ datasets with less than $10$ classes used for evaluating EquiTabPFN. The datasets are extracted from the TabZilla benchmark [16] and have a number of classes no greater than $10$. Here taskId is the OpenML ID of the task, Classes indicates the number of classes, Features the number of covariates and Samples the number of samples in each dataset.
taskId name Classes Features Samples
5 arrhythmia 12 279 360
7 audiology 23 69 180
41 soybean 19 35 545
3022 vowel 11 12 792
3481 isolet 26 617 6237
3567 collins 15 21 400
3952 chess 18 6 22444
9956 one-hundred-plants-texture 100 64 1279
125922 texture 11 40 4400
146032 primary-tumor 20 17 271
Table 2: List of datasets used for evaluating EquiTabPFN on unseen number of classes. Datasets are extracted from theTabZilla benchmark [16] and have a number of classes greater than $10$ (ranging from $11$ to $100$). Here taskId is the OpenML ID of the task, Classes indicates the number of classes, Features the number of covariates and Samples the number of samples in each dataset.
Appendix C Additional experiment results
C.1 Binary classification decision boundary
@@
C.2 Critical difference diagrams and performance metric tables
We show the critical diagram using Autorank implementation [8] in Figure 8 for datasets with more than 10 classes and Figure 7 for datasets with less than 10 classes. Critical diagrams show the average rank of each method (lower is better) and use a horizontal bar to show the methods statistically tied. We set the confidence level to $0.05$ and use default hyper-parameters while forcing non-parametric mode to ensure stability.
We also give aggregate results for datasets with more than 10 classes in Table 3 and less than 10 classes in Table 4.
Figure 7: Critical diagram on the 76 real-world datasets with less than $10$ classes from Table 1.
Figure 8: Critical diagram on the 10 real-world datasets with more than $10$ classes from Table 2.
[…]
model
EquiTabPFN 27.9 77.0 +/- 1.8 95.2 +/- 0.7 75.0 +/- 2.0 0.4
XGBoost 24.8 80.7 +/- 1.6 95.3 +/- 0.7 79.1 +/- 1.8 12.6
rtdl(ResNet) 21.2 80.9 +/- 1.7 91.1 +/- 1.1 77.9 +/- 1.9 20.5
LinearModel 16.8 65.8 +/- 2.5 92.6 +/- 0.8 63.1 +/- 2.7 0.3
RandomForest 9.4 63.3 +/- 1.5 91.6 +/- 0.7 58.4 +/- 1.6 0.8
TabPFNv2 6.0 74.5 +/- 2.6 94.4 +/- 0.9 73.6 +/- 2.6 2.8
TabPFNv2 3.9 67.3 +/- 2.5 92.7 +/- 0.9 64.6 +/- 2.7 2.8
KNN 0.0 62.9 +/- 1.9 90.3 +/- 1.0 58.4 +/- 2.1 1.6
DecisionTree -4.2 51.8 +/- 1.8 80.7 +/- 1.0 47.4 +/- 1.7 0.7
TabNet -12.5 51.2 +/- 2.8 77.5 +/- 1.7 48.7 +/- 2.9 43.4
STG -19.3 46.1 +/- 1.9 78.7 +/- 1.1 39.5 +/- 1.9 35.2
rtdl(MLP) -20.7 53.9 +/- 2.5 74.1 +/- 1.8 44.7 +/- 2.9 13.9
MLP -22.6 51.0 +/- 2.1 73.4 +/- 1.6 41.6 +/- 2.3 15.5
VIME -27.4 49.2 +/- 2.6 70.5 +/- 1.9 41.3 +/- 3.0 39.2
Table 3: Aggregate accuracy, AUC, F1 and runtime for all methods on datasets with more than 10 classes. Results are ordered by the median relative accuracy improvement w.r.t KNN over the 10 datasets after averaging over the 10 different splits. Mean accuracies, AUC and F1 score are averaged over all splits and datasets. Numbers after the symbol +/+/-+ / - refer to the standard error of the mean over all splits and datasets.
[…]
model
TabPFNv2 9.1 85.7 +/- 0.5 90.3 +/- 0.5 85.5 +/- 0.6 0.2
EquiTabPFN 6.3 83.7 +/- 0.6 88.9 +/- 0.5 83.3 +/- 0.6 0.1
XGBoost 5.2 82.7 +/- 0.6 88.1 +/- 0.5 82.5 +/- 0.6 0.4
RandomForest 4.6 79.7 +/- 0.5 86.7 +/- 0.5 78.9 +/- 0.6 0.2
TabPFNv2 4.3 80.8 +/- 0.6 87.1 +/- 0.6 80.1 +/- 0.6 0.2
rtdl(ResNet) 3.6 80.2 +/- 0.6 85.7 +/- 0.5 79.4 +/- 0.7 6.3
LinearModel 1.5 77.1 +/- 0.7 82.4 +/- 0.6 76.2 +/- 0.7 0.0
DecisionTree 0.6 76.2 +/- 0.6 79.8 +/- 0.5 75.1 +/- 0.6 0.0
MLP 0.3 73.2 +/- 0.7 76.7 +/- 0.7 71.0 +/- 0.8 6.7
rtdl(MLP) 0.3 73.4 +/- 0.8 77.3 +/- 0.7 71.1 +/- 0.9 4.8
KNN -0.0 73.8 +/- 0.6 79.5 +/- 0.6 73.0 +/- 0.7 0.0
STG -5.0 66.5 +/- 0.7 70.6 +/- 0.5 63.7 +/- 0.8 11.0
TabNet -6.4 68.5 +/- 0.6 73.5 +/- 0.5 67.5 +/- 0.7 17.2
VIME -8.1 63.5 +/- 0.7 69.7 +/- 0.7 60.8 +/- 0.8 10.5
Table 4: Aggregate accuracy, AUC, F1 and runtime for all methods on datasets with less than 10 classes. Results are ordered by the median relative accuracy improvement w.r.t KNN over the 10 datasets after averaging over the 10 different splits. Mean accuracies, AUC and F1 score are averaged over all splits and datasets. Numbers after the symbol +/+/-+ / - refer to the standard error of the mean over all splits and datasets.